Q:-If the distances between the origin and centers of the three circles x2+y2 - 2λr x -a2=0 ; r=1,2,3 are in G.P then the lengths of tangents to the circles from any point on the circle x2+y2=a2 are in__________


1.A.P


2.G.P


3.H.P


4.A.G.P


5.None of these

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Tapasranjan

x2+y2 - 2λr x -a2=0

(x-λr)2 +y2 =a2 + λr2

 center of circle( λr,0)

   so distance of center from the origin. = λr

   so λ1 2 ,λare in GP

now let a general point (acos θ ,asinθ) on circle x2+y2=a2    

so length of tangent from this point to the given circle

       = [(acos θ)2+(asinθ)2 - 2λacos θ -a2]1/2

       =[ - 2λacos θ ]1/2

 so given λ1 2 ,λare in GP

   so   -1 acos θ,-2λ2 acos θ,-3 acos θ  will also be in GP

         [ - 2λ1 acos θ ]1/2 , [ - 2λ2 acos θ ]1/2  , [ - 2λ3 acos θ ]1/2  will also be in GP

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin